# Error Correction Code Transformer: From Non-Unified to Unified

### v3

@@ Abstract
Channel coding is vital for reliable data transmission in modern wireless systems, and its significance will increase with the emergence of sixth-generation (6G) networks, which will need to support various error correction codes. However, traditional decoders were typically designed as fixed hardware circuits tailored to specific decoding algorithms, leading to inefficiencies and limited flexibility. To address these challenges, this paper proposes a unified, code-agnostic Transformer-based decoding architecture capable of handling multiple linear block codes, including Polar, Low-Density Parity-Check (LDPC), and Bose-Chaudhuri-Hocquenghem (BCH), within a single framework. To achieve this, standardized units are employed to harmonize parameters across different code types, while the redesigned unified attention module compresses the structural information of various codewords. Additionally, a sparse mask, derived from the sparsity of the parity-check matrix, is introduced to enhance the model’s ability to capture inherent constraints between information and parity-check bits, resulting in improved decoding accuracy and robustness. Extensive experimental results demonstrate that the proposed unified Transformer-based decoder not only outperforms existing methods but also provides a flexible, efficient, and high-performance solution for next-generation wireless communication systems.
@@ 1 Introduction
With the continuous evolution of wireless communication technologies, the imminent arrival of sixth-generation (6G) wireless communications heralds a paradigm shift towards ultra-reliable, low-latency communications and notably enhanced data rates  [1, 2, 3]. To fulfill the ambitious performance requirements of 6G networks, probabilistic coding schemes are anticipated to play a crucial role. In particular, linear block codes such as Polar codes  [4] and Low-Density Parity-Check (LDPC) codes  [5] are expected to maintain their competitive edge established in fifth-generation (5G) wireless communications. These codes have also emerged as promising candidates for efficient error correction in 6G  [6]. Additionally, another algebraic linear block code, known as the Bose-Chaudhuri-Hocquenghem (BCH) code  [7], is often used as an outer code to enhance decoding performance.
@@ 1.1 Prior Works
Looking at 5G and earlier wireless communication systems, linear block code decoders were typically designed as fixed hardware circuits tailored to specific decoding algorithms. Looking ahead to 6G, this trend is likely to persist  [8]. For instance, the Successive Cancellation (SC) decoding algorithm  [4] for Polar codes uses a sequential method that traverses a decision tree to obtain the decoding result. This approach simplifies the hardware implementation of Polar codes by pruning binary trees  [9, 10, 11]. In contrast, the Sum-Product (SP) decoding algorithm  [12] for LDPC codes uses iterative message passing, allowing for parallel data processing. This fundamentally differs from the sequential method used for Polar codes in hardware implementation  [13]. Furthermore, BCH codes typically use algebraic decoding techniques  [14], classified as hard decoding, which differs greatly from soft decoding methods like SC and SP. Overall, these algorithms provide strong error correction capabilities. However, due to the non-unified nature of the sequential and parallel architectures of Polar and LDPC codes, and the gap between hard and soft decoding in BCH codes, unique hardware circuits must be customized and fine-tuned for each specific linear block code. This may lead to significant engineering effort and inefficient use of hardware resources.
In an effort to reduce hardware resource consumption and establish a quasi-unified decoding framework, three methods have been explored: resource sharing, Ordered-Statistics Decoding (OSD), and Artificial Intelligence (AI)-driven approaches. Resource sharing involves time-sharing storage and computing units. Ordered-Statistics Decoding leverages the reliability ordering of received symbols to efficiently decode various linear block codes by systematically evaluating the most likely codewords. Meanwhile, AI-driven methods offer a convenient way to unfold traditional algorithms like Belief Propagation (BP).
@@
First, resource-sharing methods, such as those in  [15] and  [16], endeavor to integrate decoders for Polar and LDPC codes into a shared architecture based on traditional decoding algorithms. Both approaches aim to conserve hardware resources by temporally multiplexing the computational and storage units shared by the decoder through reconfiguration. However, due to the distinct decoding algorithms for each code type, achieving a truly unified architecture with such decoders is generally challenging.
@@
Second, ordered-statistics decoding methods, initially explored in  [17, 18, 19], provide a near-Maximum Likelihood (ML) decoding approach for linear block codes by exploiting the reliability ordering of received symbols. The algorithm operates by sorting the received signal positions based on their reliability, permuting the code’s generator matrix accordingly, and systematically evaluating the most likely codewords through a limited search. OSD delivers near-ML performance with significantly less complexity than exhaustive ML decoding, making it well-suited for short to moderate-length codes. However, its drawback lies in the computational complexity, which grows exponentially with the search order (e.g., order-$l$). For instance, an order-4 OSD may require evaluating thousands of test codewords, potentially leading to increased latency and resource demands.
Third, in recent years, AI-driven decoders have gained attention as an alternative. While model-based AI methods leverage traditional decoder algorithms and rely on neural networks to simulate belief propagation, their performance improvements are often limited by the constraints of conventional models  [20],  [21]. In contrast, model-free approaches offer a more flexible and scalable solution. Unlike model-based methods, which depend on predefined mathematical models, model-free approaches assume no prior knowledge of the encoder or decoder. Instead, they rely on large amounts of labeled data to learn the decoding process directly, bypassing the limitations of conventional algorithms  [22].
@@
For instance, inspired by the success of Transformers in Natural Language Processing (NLP)  [23], a model-free neural decoder called the Error Correction Code Transformer (ECCT) was proposed  [24]. ECCT employs the Transformer’s self-attention mechanism, where multiple attention heads operate in parallel to process sequential data and capture long-range dependencies, effectively linking input Log-Likelihood Ratios (LLRs) to the decoder’s output. The inherent connection between Transformers and decoders lies in their shared ability to translate sequences—whether from one language to another or from LLRs back to the original information bits. However, ECCT’s flexibility comes at a cost, requiring frequent parameter adjustments for different code types, lengths, and rates, which increases training complexity and hardware overhead, hindering a unified decoding architecture. Building on this,  [25] introduced a Foundation Model for Error Correction Codes (FECCT) to improve ECCT’s generalization to unseen codewords. Yet, FECCT currently faces the challenge of high computational complexity in its self-attention mechanism, further complicating its practical implementation.
@@ 1.2 Contributions
Thus, to reduce engineering efforts, training costs, and hardware consumption, we propose a unified, code-agnostic Transformer-based decoding architecture capable of handling multiple linear block codes, including Polar, LDPC, and BCH, within a single framework. To achieve this, we designed new standardized units and a unified attention module. Additionally, a sparse masking mechanism was developed to enhance decoding performance and stability  [^1] Simulation codes will be provided to reproduce the results in this paper: http://oa.ee.tsinghua.edu.cn/dailinglong/publications/publications.html. Specifically, the main contributions of this paper are summarized as follows.
@@
    Unified Decoding Architecture: We introduce a unified, code-agnostic Transformer decoding architecture, capable of seamlessly integrating various linear block codes. This architecture features standardized units to align various code lengths and, for the first time, designs a unified attention module that learns to compress the structural information of all codewords, thereby achieving a code-agnostic decoder. This architecture obviates the necessity for distinct hardware circuits for different decoding algorithms, reducing engineering efforts and hardware resource inefficiency.
@@
    Standardized Units: Standardized units are used to align parameters such as code length and code rate across different code types. Specifically, after receiving the channel output, we pad different codewords and their corresponding syndromes with zeros up to the maximum length. Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner, thereby reducing training overhead.
@@
    Unified Attention: We redesigned the self-attention module to create a unified attention mechanism that allows different attention heads to share attention scores. This approach contrasts with traditional self-attention mechanisms, where each head independently computes its scores using separate learned weights. The unified attention compresses the structural information of different codewords through learning and shares parameters across various heads in the Transformer, resulting in a more cohesive decoding architecture.
@@
    Performance Enhancement: By introducing a sparse mask, which leverages the sparsity of the parity-check matrix, we have significantly enhanced the model’s ability to capture inherent constraints between information and parity-check bits. This has led to a substantial improvement in decoding performance and the achievement of a state-of-the-art design.
  5. 5.
    High Computational Efficiency: The unified attention module in our proposed Transformer-based decoding architecture demonstrates linear computational complexity by eliminating the need to project the input into higher dimensions. This results in a significant reduction in complexity from the traditional $\mathcal{O}(N^{2})$ to a more efficient $\mathcal{O}(N)$, where $N$ represents the sequence length. Additionally, by employing our sparse mask, we further achieve an $86\%$ reduction in computational complexity. This efficiency gain enables faster processing and lower power consumption, making the architecture suitable for real-time applications in next-generation wireless communication systems.
The rest of this paper is organized as follows. Section II provides background knowledge, including the vanilla Transformers, channel encoders, as well as pre- and post-processing involved to address the issue of overfitting. Section III details our proposed unified code-agnostic Transformer. Section IV presents the results of training and inference. Section V analyzes the impact of the proposed optimization algorithms. Finally, Section VI concludes.
@@ 2.1 Vanilla Transformers
The vanilla Transformer, a sequence-to-sequence model, comprises an encoder and decoder, each with $L$ identical blocks. Each block integrates multi-head self-attention (MHA), a position-wise feed-forward network (FFN), residual connections  [26] to mitigate gradient issues, and layer normalization  [27] for stability and faster convergence, as shown in Figure 1.
@@
$\textbf{{X}}_{e}=\text{Embedding}(\textbf{{x}})=\text{diag}(\textbf{{x}})\cdot\textbf{{W}}$ (1)
@@
$\text{Attn}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}}{\sqrt{d_{k}}}\right)\cdot\textbf{{V}}$ (2)
@@
$\text{MHA}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Concat}\left(\text{head}_{1},\cdots,\text{head}_{H}\right)\cdot\textbf{{W}}^{\textit{O}}$ (3)
where $\text{head}_{i}=\text{Attn}(\textbf{{X}}_{Q}\textbf{{W}}_{i}^{Q},\textbf{{X}}_{K}\textbf{{W}}_{i}^{K},\textbf{{X}}_{V}\textbf{{W}}_{i}^{V})$, with $\textbf{{X}}_{Q},\textbf{{X}}_{K},\textbf{{X}}_{V}\in\mathbb{R}^{N\times d_{k}}$ set to $\textbf{{X}}_{e}$ for self-attention, using trainable weights $\textbf{{W}}_{i}^{Q},\textbf{{W}}_{i}^{K},\textbf{{W}}_{i}^{V}\in\mathbb{R}^{d_{k}\times d_{k}}$ and $\textbf{{W}}^{O}\in\mathbb{R}^{Hd_{k}\times Hd_{k}}$, as depicted in Figure 2.
@@
$\text{FFN}\left(\textbf{{X}}\right)=\textbf{{W}}_{2}\cdot\text{ReLU}\left(\textbf{{W}}_{1}\cdot\textbf{{X}}+\textbf{{b}}_{1}\right)+\textbf{{b}}_{2}$ (4)
@@
$\textbf{{X}}_{2}$ $=\text{FFN}\left(\text{LayerNorm}\left(\textbf{{X}}_{1}\right)\right)+\textbf{{X}}_{1}$
@@ 2.2 Forward Error Correction Codes
The encoder maps a message vector $\textbf{{m}}\in\{0,1\}^{k}$ to a codeword $\textbf{{x}}\in\{0,1\}^{n}$ via $\textbf{{x}}=\textbf{{m}}\cdot\textbf{G}$, where $\textbf{H}\cdot\textbf{{x}}=\textbf{0}$. This codeword is modulated using Binary Phase Shift Keying (BPSK), defined as $\text{BPSK}(x)=1-2x$, yielding $\textbf{{x}}_{s}$, which is transmitted over Binary-Input Discrete Memoryless Channels (B-DMCs, e.g., AWGN), producing $\textbf{{y}}=\textbf{{x}}_{s}+\textbf{{n}}$, with $\textbf{{n}}\sim\mathcal{N}(0,\sigma^{2})$.
@@
The decoder aims to recover m from y, as depicted in Figure 3. Our work focuses on designing and training a parameterized decoding function $f_{\theta}$. To develop a model-free Transformer-based decoder, we leverage pre- and post-processing techniques from  [28], preserving BER and Mean Squared Error (MSE) performance (see Figure 4). Training with an all-zero codeword mitigates overfitting due to exponential codeword space growth. And the B-DMCs is modeled by (6) which is an equivalent statistical mode that differs from the true physical one  [29].
$\textbf{{y}}=\textbf{{x}}_{s}\cdot\textbf{{z}}$ (6)
@@
$\widetilde{\textbf{{y}}}=\left[\lvert\textbf{{y}}\rvert,\text{s}\left(\textbf{{y}}\right)\right]$ (7)
@@ 3 Proposed Unified Error Correction Code Transformer
In this section, we present the complete architecture and training process of the proposed Unified Error Correction Code Transformer (UECCT).

### v4

# Unified Error Correction Code Transformer with Low Complexity

@@ Abstract
Channel coding is vital for reliable sixth-generation (6G) data transmission, employing diverse error correction codes for various application scenarios. Traditional decoders require dedicated hardware for each code, leading to high hardware costs. Recently, artificial intelligence (AI)-driven approaches, such as the error correction code Transformer (ECCT) and its enhanced version, the foundation error correction code Transformer (FECCT), have been proposed to reduce the hardware cost by leveraging the Transformer to decode multiple codes. However, their excessively high computational complexity of $\mathcal{O}(N^{2})$ due to the self-attention mechanism in the Transformer limits scalability, where $N$ represents the sequence length. To reduce computational complexity, we propose a unified Transformer-based decoder that handles multiple linear block codes within a single framework. Specifically, a standardized unit is employed to align code length and code rate across different code types, while a redesigned low-rank unified attention module, with computational complexity of $\mathcal{O}(N)$, is shared across various heads in the Transformer. Additionally, a sparse mask, derived from the parity-check matrix’s sparsity, is introduced to enhance the decoder’s ability to capture inherent constraints between information and parity-check bits, improving decoding accuracy and further reducing computational complexity by $86\%$. Extensive experimental results demonstrate that the proposed unified Transformer-based decoder outperforms existing methods and provides a high-performance, low-complexity solution for next-generation wireless communication systems.
@@ 1 Introduction
With the continuous evolution of wireless communication technologies, the imminent arrival of sixth-generation (6G) wireless communication will mark a paradigm shift toward ultra-reliable communication [1, 2, 3, 4]. To meet the stringent performance demands of 6G, channel codes with strong error correction capabilities are expected to play a pivotal role. Specifically, low-density parity-check (LDPC) codes [5] and Polar codes [6] have proven effective in fifth-generation (5G) wireless communications. Due to their excellent performance, 3GPP has just made the decision to continue to use LDPC and Polar codes for future 6G in October 2025 [7]. Furthermore, algebraic geometry codes, such as Bose-Chaudhuri-Hocquenghem (BCH) codes [8], are commonly employed as outer codes in concatenated schemes with probabilistic codes to enhance decoding performance.
@@ 1.1 Prior Works
In wireless communication systems, multiple coding schemes are supported, such as LDPC and Polar codes in 5G [9], each requiring customized hardware for their decoding algorithms. For instance, LDPC codes employ iterative message passing for parallel data processing [10], while Polar codes rely on the successive cancellation (SC) decoding algorithm [6], which follows a sequential approach by traversing a decision tree, often simplified through binary tree pruning [11, 12, 13, 14]. Both LDPC and Polar codes utilize soft decoding methods, leveraging probabilistic information to achieve robust error correction. In contrast, BCH codes, another key coding scheme, employ algebraic decoding techniques classified as hard decoding [15], which rely on mathematical structures to correct errors.The gap between the hard decoding of BCH codes and the soft decoding of LDPC and Polar codes, combined with the distinct parallel and sequential architectures of LDPC and Polar codes, necessitates customized hardware circuits for each code type, resulting in significant hardware resource cost.
@@
To reduce hardware resource cost, three approaches have been investigated for decoding multiple codecs with a single hardware framework: resource sharing, ordered statistical decoding, and artificial intelligence (AI)-driven methods. First, resource-sharing methods, such as those in [16, 17], integrate LDPC and Polar code decoders into a single architecture using traditional decoding algorithms. These methods conserve hardware resources by multiplexing shared computational and storage units through reconfiguration. However, due to the distinct decoding algorithms for each code type, only partial resource sharing is achieved, limiting hardware resource reduction.
@@
Second, ordered-statistics decoding methods, as explored in [18, 19, 20], provide near-maximum likelihood (ML) decoding for linear block codes by ordering received symbols by reliability, enabling hardware reuse and reducing resource costs. Ordered-statistics decoding sorts signal positions by reliability, permutes the generator matrix of the code in a consistent manner across different linear block codes, and evaluates the most likely codewords through a limited search. This uniform approach to generator matrix processing enables hardware reuse across various linear block codes, reducing resource cost compared to code-specific decoding architectures. However, the computational complexity of ordered-statistics decoding increases exponentially with code length and search order (e.g., order-$l$). For example, an order-4 ordered-statistics decoding may require evaluating thousands of test codewords, increasing decoding latency.
@@
Third, in recent years, AI-driven decoders have gained attention as an alternative to reduce hardware resource cost. For example, inspired by the success of Transformers in natural language processing [21], the error correction code Transformer (ECCT) was introduced to reduce hardware resource and enhance decoding performance for short codes [22]. The inherent connection between Transformers and decoders lies in their shared capability for sequence-to-sequence transformations, mapping input sequences—e.g., from one language to another in natural language processing or from log-likelihood ratios (LLRs) to information bits in channel decoding—into output sequences. ECCT leverages the Transformer’s self-attention mechanism, where multiple attention heads operate in parallel to process sequential data and capture long-range dependencies, effectively modeling the constrained relationships among the LLRs of decoder inputs. Building on ECCT, the foundation error correction code Transformer (FECCT) enhances generalization to unseen codewords by integrating the distance matrix, derived from the parity-check matrix $\mathbf{H}$, into the self-attention mechanism [23]. However, ECCT and FECCT rely on the Transformer’s powerful self-attention mechanism, which has a computational complexity of $\mathcal{O}(N^{2})$, where $N$ represents the sequence length. Their high computational complexity makes them difficult to scale to long codes.
@@ 1.2 Contributions
To reduce computational complexity, we propose a unified Transformer-based decoder capable of handling multiple linear block codes, including LDPC, Polar, and BCH, within a single framework. To achieve this, we design a standardized unit and a low-rank unified attention module that shares attention scores across various heads in the Transformer. Additionally, we develop a sparse masking mechanism to enhance decoding performance and further reduce computational complexity [^1] Simulation codes will be provided to reproduce the results in this paper: http://oa.ee.tsinghua.edu.cn/dailinglong/publications/publications.html.. Specifically, the main contributions of this paper are summarized as follows.
@@
    We introduce a unified Transformer-based decoder capable of seamlessly integrating various linear block codes. This architecture incorporates a standardized unit to align code lengths, a unified attention module that compresses the structural information of all codewords, and a sparse mask leveraging the parity-check matrix’s sparsity to enhance decoding accuracy. These components achieve a code-agnostic decoder, eliminate the need for distinct hardware circuits for different decoding algorithms, and reduce computational complexity.
@@
    A standardized unit aligns parameters, such as code length and code rate, across different code types. Specifically, it pads codewords and their corresponding syndromes with zeros to a maximum length after receiving the channel output. This enables the Transformer-based decoder to handle any codeword within the maximum length, facilitating unified joint training.
@@
    We redesigned the self-attention module to create a unified attention mechanism that allows different attention heads to share attention scores. This approach contrasts with traditional self-attention mechanisms, where each head independently computes its scores using separate learned weights. This unified attention compresses structural information of different codewords through learning and shares parameters across various heads in the Transformer. It eliminates the need to project inputs into higher dimensions, reducing computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$, where $N$ represents the sequence length. This results in a more efficient and unified decoding architecture.
@@
    By introducing a sparse mask, which leverages the sparsity of the parity-check matrix, we have significantly enhanced the model’s ability to capture inherent constraints between information and parity-check bits. This has led to a substantial improvement in decoding performance and the achievement of a state-of-the-art design. Additionally, this sparse mask further reduces computational complexity by $86\%$, enabling efficient training for long codes.
The rest of this paper is organized as follows. Section II provides background knowledge, including the vanilla Transformers, channel encoders, as well as pre- and post-processing involved to address the issue of overfitting. Section III details our proposed unified Transformer-based decoder. Section IV presents the results of training and inference. Section V analyzes the impact of the proposed optimization algorithms. Finally, Section VI concludes.
@@ 2.1 Vanilla Transformers
The vanilla Transformer, a sequence-to-sequence model, comprises an encoder and decoder, each with $L$ identical blocks. Each block integrates multi-head self-attention (MHA), a position-wise feed-forward network (FFN), residual connections to mitigate gradient issues [24], and layer normalization for stability and faster convergence [25], as shown in Figure 1.
@@
$\textbf{{X}}_{e}=\text{Embedding}(\textbf{{x}})=\text{diag}(\textbf{{x}})\cdot\textbf{{W}},$ (1)
@@
$\text{Attn}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}}{\sqrt{d_{k}}}\right)\cdot\textbf{{V}},$ (2)
@@
$\text{MHA}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Concat}\left(\text{head}_{1},\cdots,\text{head}_{H}\right)\cdot\textbf{{W}}^{\textit{O}},$ (3)
where $\text{head}_{i}=\text{Attn}(\textbf{{X}}_{Q}\textbf{{W}}_{i}^{Q},\textbf{{X}}_{K}\textbf{{W}}_{i}^{K},\textbf{{X}}_{V}\textbf{{W}}_{i}^{V})$, with $\textbf{{X}}_{Q},\textbf{{X}}_{K},\\ \textbf{{X}}_{V}\in\mathbb{R}^{N\times d_{k}}$ set to $\textbf{{X}}_{e}$ for self-attention, using trainable weights $\textbf{{W}}_{i}^{Q},\textbf{{W}}_{i}^{K},\textbf{{W}}_{i}^{V}\in\mathbb{R}^{d_{k}\times d_{k}}$ and $\textbf{{W}}^{O}\in\mathbb{R}^{Hd_{k}\times Hd_{k}}$, as depicted in Figure 2.
@@
$\text{FFN}\left(\textbf{{X}}\right)=\textbf{{W}}_{2}\cdot\text{ReLU}\left(\textbf{{W}}_{1}\cdot\textbf{{X}}+\textbf{{b}}_{1}\right)+\textbf{{b}}_{2},$ (4)
@@
$\textbf{{X}}_{2}$ $=\text{FFN}\left(\text{LayerNorm}\left(\textbf{{X}}_{1}\right)\right)+\textbf{{X}}_{1}.$
@@ 2.2 Forward Error Correction Codes
The encoder maps a message vector $\textbf{{m}}\in\{0,1\}^{k}$ to a codeword $\textbf{{x}}\in\{0,1\}^{n}$ via $\textbf{{x}}=\textbf{{m}}\cdot\textbf{G}$, where $\textbf{H}\cdot\textbf{{x}}=\textbf{0}$. This codeword is modulated using binary phase shift keying (BPSK), defined as $\text{BPSK}(x)=1-2x$, yielding $\textbf{{x}}_{s}$, which is transmitted over binary-input discrete memoryless channels (B-DMCs, e.g., additive white Gaussian noise (AWGN) channels), producing $\textbf{{y}}=\textbf{{x}}_{s}+\textbf{{n}}$, with $\textbf{{n}}\sim\mathcal{N}(0,\sigma^{2})$.
@@
The decoder aims to recover m from y, as depicted in Figure 3. Our work focuses on designing and training a parameterized decoding function $f_{\theta}$. To develop a data-driven Transformer-based decoder, we leverage pre- and post-processing techniques from [26], preserving BER and mean squared error (MSE) performance (see Figure 4). Training with an all-zero codeword mitigates overfitting due to exponential codeword space growth. And the B-DMCs is modeled by (6) which is an equivalent statistical mode that differs from the true physical one [27].
$\textbf{{y}}=\textbf{{x}}_{s}\cdot\textbf{{z}},$ (6)
@@
$\widetilde{\textbf{{y}}}=\left[\lvert\textbf{{y}}\rvert,\text{s}\left(\textbf{{y}}\right)\right],$ (7)
@@ 3 Proposed Unified Error Correction Code Transformer
In this section, we present the detailed architecture and training process of the proposed unified error correction code Transformer (UECCT).